The length of a rectangle is 7ft more than twice the width and the area of the rectangle is 99ft squared .find the dimensions of the rectangle
Found 2 solutions by Alan3354, addingup:
Answer by Alan3354(69443)

The length of a rectangle is 7ft more than twice the width and the area of the rectangle is 99ft squared .find the dimensions of the rectangle
---------------
L = 2W + 7
L*W = 99

Answer by addingup(3677)

L= 2W+7
L*W= 99
(2W+7)*W = 99
2W^2+7W = 99
W^2+7/2W = 99/2
Take 1/2 the coefficient and square it. The coefficient is 7/2 and (7/2)/2
7/2*1/2 = 7/4 Now we square it:(7/4)^2 = 49/16 add this to both sides:
W^2+7/2+49/16 = 99/2+49/16
W^2+7/2+49/16 = 841/16
Perfect square on the left (x+7/2)*(x+7/2), multiply First Outer Inner Last and you get x^2+7/2x+49/16
(x+7/2)*(x+7/2) can be written as (x+7/2)^2:
(x+7/2)^2 = 841/16
x+7/8 = sqrt 841/16
x = -9
or
x = 11/2 = 5.5 fraction or decimal, depends how you need it.
We know we're looking for a positive number, so let's try 5.5 in the formula for the area:
(2*5.5)+7*5.5 = 99
18*5.5 = 99 Correct
:
John